Due to the intervention on SDB2 performed 2 weeks ago, the power sent to the B1p and B5 quadrants has been increased.
One can see on Fig.1 and Fig.2 that the H and V signals are saturating on the B1p and B5 quadrants when the CITF is flashing.
I closed the shutters of the B1p quadrants at 16h09 utc and the shutters of the B5 quadrants at 16h18 utc.
In order to keep these quadrants safe, we must keep their shutters closed during the alignment of the CITF. They can be opened once DRMI is locked.